OpenStudy (anonymous):

What is 5/5^1/3 in radical form? Help would be greatly appreciated.

OpenStudy (anonymous):

is it like this \[\frac{ 5 }{ 5^{\frac{ 1 }{ 3 }} }\text{ or }\left( \frac{ 5 }{5 } \right)^{\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

It's like the first one ^^

OpenStudy (shamil98):

That is the same as \[\large \frac{ 5 }{ \sqrt[3]{5} }\]

OpenStudy (shamil98):

basically the formula is. \[\large \sqrt[n]{x^m} = x^{\frac{ m }{ n }}\]

OpenStudy (anonymous):

So would the answer me 1/Cubed root of square root of 25? An answer would help out D;

OpenStudy (anonymous):

\[\frac{ x^b }{ x^a }=x^{(b-a)}\]

OpenStudy (anonymous):

I don't think that formula helped me get the answer :c

OpenStudy (anonymous):

\[\frac{ 5^1 }{ 5^{\frac{ 2 }{ 3 } }}=5^{1-\frac{ 1 }{ 3 }}=\]

OpenStudy (anonymous):

then use @shamil98's formula

OpenStudy (anonymous):

I got Cubed root of square root of 25! I think thats it c: Thanks a bunch.